Please help me: A rectangular garden has a perimeter of 140m and an area of 1200m^2. set up a quadratic equation and solve it agebraically to find the dimensions of the garden.
Answer by checkley77(12844)

2L+2W=140
LW=1,200
L=1,200/W
2(1,200/W)+2W=140
2,400/W+2W=140
(2,400+2W*W)/W=140
2,400+2W^2=140W
2W^2-140W+2,400=0
2(W^2-70W+1,200)=0
2(W-30)(W-40)=0
W-30=0
W=30 L=40 ANS.
W-40=0
W=40 L=30 ANS.
PROOF:
30*40=1,200
2*30+2*40=140
60+80=140
140=140